# RSAdapter: Adapting Multimodal Models for Remote Sensing Visual Question Answering

### v1

@@ IV-D5 Different Position of RSAdapters
IV-D6 Different Number of Transformer Layers
Finally, to verify the effectiveness of the model under varying computational resources, we conduct experiments with different numbers of transformer layers, as shown in Table X. The results demonstrate that even with only 6 transformer layers, we can still achieve superior performance compared to all previous models. This reaffirms the effectiveness of our method.
TABLE X: Effect of different number of transformer layers on the LR test set. Results are reported as percentage (%). Best results are highlighted in bold.
@@ IV-D6 Different Number of Transformer Layers
IV-E Visulization
@@
IV-F Complexity Analysis
TABLE XI: Comparison of complexity with RSVQA [14] on LR dataset.
Methods Param (M) Tunable Param (M) Training Time (s) Testing Time (s)
RSVQA [14] 85.7 5.7 154 28
ours 120 8.4 159 16
Here, we evaluate the complexity of RSAdapter. We assume that the model’s feature dimension is represented as $d$, and the bottleneck size of RSAdapter is denoted as $d^{\prime}$. Consequently, the tunable parameters for each transformer layer during the training phase can be calculated as $2(dd^{\prime}+2(d+d^{\prime}))$. During inference, the number of extra parameters is simplified to $2dd^{\prime}$.
Furthermore, we conduct a series of complexity comparisons with RSVQA [14], and the results are presented in Table XI. The training and testing times represent the average time for one epoch, and the experiment was conducted on an Nvidia RTX A6000 GPU. From the results, we can infer that our method requires a similar amount of time for training one epoch compared to the RSVQA model, even though it has slightly more tunable parameters. However, our method achieves nearly twice the inference speed of RSVQA. Notably, SHRNet’s training time is 3.8 times longer than RSVQA, and its testing time is 3.3 times longer than RSVQA based on [18]. This strongly suggests that our model provides a substantial improvement in runtime efficiency compared to SHRNet. In terms of parameter comparison, SHRNet has 105.56 million training parameters, whereas we achieve better results on all three datasets with only 10% of SHRNet’s parameter count. These experiments demonstrate the runtime and parameter efficiency of our method.
@@ V Conclusions
In future work, we aim to achieve efficient fine-tuning by pre-training multimodal models on a large number of remote sensing images, thereby avoiding the loss associated with transferring from natural images to remote sensing images. Additionally, we hope to discover better methods to further reduce the size of the training parameters without sacrificing performance as much as possible.

### v2

@@ IV-D6 Different Position of RSAdapters
IV-D7 Different Number of Transformer Layers
Finally, to verify the effectiveness of the model under varying computational resources, we conduct experiments with different numbers of transformer layers, as shown in Table XI. The results demonstrate that even with only 6 transformer layers, we can still achieve superior performance compared to all previous models. This reaffirms the effectiveness of our method.
TABLE XI: Effect of different number of transformer layers on the LR test set. Results are reported as percentage (%). Best results are highlighted in bold.
@@ IV-D7 Different Number of Transformer Layers
IV-E Language Bias
For current RS VQA models, a significant concern is language bias. To verify whether our approach learns strong language biases, we conduct experiments on the RSIVQA dataset. Firstly, we define three scenarios: in Scenario 1, the input during training consists of original image-question pairs; in Scenario 2, the input remains the same, but during testing, the input image is replaced with a random test image, implying a high probability of mismatch between the image and the question. In the final Scenario 3, we train the model using questions and random images to assess its generalization ability. Table XII presents the experimental results. Compared to Full finetune, RSAdapter exhibits more decrease in accuracy when the test images are replaced. This indicates that our method learns more features from the images and less language bias. This result is also evident from the results of Scenario 3. The accuracy of our method is lower than that of Full finetune, suggesting that in the same circumstances, Full finetune relies more on questions and learns more language bias. Therefore, compared to Full finetune, our model exhibits stronger generalization and learns less language bias.
TABLE XII: Comparison of the performance of Full finetune and RSAdapter on the RSIVQA dataset in three different scenarios. All values are reported as percentage (%).
Methods Scenario Types Overall Accuracy
Yes/No Number Others
Full finetune 1 97.94 64.24 90.10 87.27
2 78.16 47.53 4.34 59.61
3 84.04 49.01 7.80 63.91
RSAdapter 1 97.90 62.64 92.47 87.10
2 70.73 42.90 5.22 54.08
3 83.69 48.69 7.73 63.60
Note: In Scenario 1, the input consists of the original images and questions from the dataset. In Scenario 2, the input is the same as Scenario 1, but random images are used during testing. In Scenario 3, the input consists of questions and random images.
IV-F Visulization
@@
Figure 6: Typical remote sensing visual question answering examples on LR (a), HR (b) and RSIVQA (c) datasets. From top to bottom, the figure includes the original image, attention map over image, and attention map over text.
@@
IV-G Complexity Analysis
TABLE XIII: Comparison of complexity with other methods on LR dataset.
Methods Param (M) Tunable Param (M) Training Time (s) Testing Time (s) Overall Accuracy(%)
RSVQA [14] 85.7 5.7 154 28.04 79.08
Bitfit [36] 113 1.3 158 15.63 84.72
Lora [32] 120 8.3 157 15.80 86.37
Adapter [29] 120 8.4 159 15.97 86.55
RSAdapter w/o rep 120 8.4 159 16.59 87.14
RSAdapter w/ rep 120 8.4 159 16.03 87.14
Here, we evaluate the complexity of RSAdapter. We assume that the model’s feature dimension is represented as $d$, and the bottleneck size of RSAdapter is denoted as $d^{\prime}$. Consequently, the tunable parameters for each transformer layer during the training phase can be calculated as $2(dd^{\prime})+2(d+d^{\prime})$. During inference, the number of extra parameters is simplified to $2dd^{\prime}$.
Furthermore, we conduct a series of complexity comparisons with RSVQA [14], and the results are presented in Table XIII. The training and testing times represent the average time for one epoch, and the experiment was conducted on an Nvidia RTX A6000 GPU. From the results, we can infer that our method requires a similar amount of time for training one epoch compared to the RSVQA model, even though it has slightly more tunable parameters. However, our method achieves nearly twice the inference speed of RSVQA. Notably, SHRNet’s training time is 3.8 times longer than RSVQA, and its testing time is 3.3 times longer than RSVQA based on [18]. This strongly suggests that our model provides a substantial improvement in runtime efficiency compared to SHRNet. In terms of parameter comparison, SHRNet has 105.56 million training parameters, whereas we achieve better results on all three datasets with only 10% of SHRNet’s parameter count. In addition, we compare our approach with other common PEFT methods. Compared to Bitfit and Lora, RSAdapter requires longer training and testing times due to the introduction of new modules by the Adapter. However, compared to Adapter, using RSAdapter results in a significant improvement in OA within the same training duration. Additionally, the model’s inference speed increased by 3.38% after applying re-parameterization techniques. These experiments demonstrate the runtime and parameter efficiency of our method.
IV-H Application in RS
IV-H1 Scale Variation
Images captured in remote sensing applications can vary significantly in scale, from large-scale satellite images to close-up aerial imagery. Addressing scale variation in image-text models is crucial for accurately understanding and interpreting remote sensing data.
IV-H2 Dataset Diversity
Remote sensing datasets often encompass diverse environments, terrains, and conditions. Ensuring that image-text models are trained on diverse datasets representative of various remote sensing scenarios is essential for robust performance across different contexts.
IV-H3 Language Biases
Language biases in RS VQA datasets can skew model performance and limit generalization, particularly in remote sensing applications. Mitigating language biases through balanced dataset curation and careful model training is essential.
IV-H4 Model Interpretability
Interpretable models are crucial in remote sensing applications, where decisions based on model outputs may have significant real-world consequences. Ensuring that image-text models provide interpretable explanations for their predictions can enhance trust and facilitate decision-making in remote sensing tasks.
@@ V Conclusions
In future work, we aim to achieve efficient fine-tuning by pre-training multimodal models on a large number of remote sensing images, thereby avoiding the loss associated with transferring from natural images to remote sensing images. Additionally, we hope to discover better methods to further reduce the size of the training parameters without sacrificing performance as much as possible. Current RS-VQA datasets are designed based on templates, limiting the assessment of current algorithms to simpler questions and their reasoning abilities. A valuable research direction would be to create more complex benchmark datasets for visual question answering in the remote sensing community.